OpenStudy (anonymous):

What are the solution of the quadratic equation? 5x-squared = 10x

OpenStudy (anonymous):

\[5x^2-10x=0\] 5x(x-2)=0 x=0, x-2=0 x=0,2

OpenStudy (anonymous):

Ty x3